# Programmable Optical Filters Based on Feed-Forward Photonic Meshes

### v2

@@ Abstract
Traditional fixed-function photonic filters offer high performance but lack flexibility, requiring separate device designs and fabrication runs for each target function. Programmable filters enable a single hardware platform to implement a wide range of spectral responses, enabling further applications including all-optical signal processing [14] and correlation spectroscopy [15]. Programmable spectral filters have been realized utilizing a number of recirculating mesh architectures [14, 16, 17, 18]. In these systems, light is routed through closed-loop paths that provide spectral selectivity via interference and resonance effects.
@@ Results
For a photonic mesh represented by the matrix elements $O_{qp}$, the transfer function to each output of the programmable spectral filter is given by Eq. 9 where $x(\Delta\omega)$ is the input spectrum to the PIC and $y_{q}(\Delta\omega)$ is the spectrum at the $q^{th}$ output. :
@@
$y_{q}(\Delta\omega)=[\sum_{p=0}^{N-1}a_{p}O_{qp}e^{-jk_{p}\Delta\omega\Delta{t}\ -2\pi\gamma}]x(\Delta\omega)$ (11)
@@
An additional dynamic, wavelength-independent phase profile may be introduced using programmable phase shifters to perturb the static phase profile generated by the array of waveguide delay lines. This dynamic phase profile may be generated either using the $\phi_{00}$, $\phi_{01}$, and $\phi_{02}$ phase shifters in the $m=0$ layer of the photonic mesh or the $\phi_{0}$, $\phi_{1}$, and $\phi_{2}$ phase shifters of the power splitting sub-circuit. Whether this phase profile is introduced prior to or following the array waveguide delay lines is of no consequence. By programming these phase shifters to introduce a linear tilt in addition to the phase shifts determined by self-configuration, the center wavelength of a preconfigured filter function may be shifted. The programmable linear phase profile is defined by the terms $e^{-j2\pi\gamma{p}}$. By adjusting the value of gamma between 0 and 1, the center wavelength of the preconfigured filter function may be tuned continously across an FSR. This principle is demonstrated for a $N=4,M=1$ spectral filter for a selection of $\gamma$ values in figure 4. The ability to modulate the center wavelength of a preconfigured filter function is a prerequisite for applications such as correlation spectroscopy used in environmental and biomedical sensing [15].
@@ Discussion
Among potential applications, this device enables tunable pump wavelength rejection in Raman spectroscopy, continuously tunable bandpass filters for use in external cavity lasers, and reprogrammable synthetic spectra generation for use in correlation spectroscopy. The ability to modulate the central wavelength of a preconfigured filter function, enabled by this device, is a prerequisite of filters used in correlation spectroscopy. Further, we have demonstrated the capability of the device to perform DWDM with inter-channel crosstalk between -25 dB and -40 dB. Unlike typical DWDM systems that operate on a static wavelength grid, mapping individual wavelengths sequentially to individual output channels, this device may dynamically shift the wavelength grid and dynamically remap any input wavelength channel to any output waveguides. In essence, this further introduces the functionality of a switching network.

### v3

@@ Abstract
Traditional fixed-function photonic filters offer high performance but lack flexibility, requiring separate device designs and fabrication runs for each target function. Programmable filters enable a single hardware platform to implement a wide range of spectral responses, enabling further applications including all-optical signal processing [14], correlation spectroscopy [15] and wavelength tunability for external cavity lasers [16]. Programmable spectral filters have been realized utilizing a number of recirculating mesh architectures [14, 17, 18, 19]. In these systems, light is routed through closed-loop paths that provide spectral selectivity via interference and resonance effects.
@@ Results
For a photonic mesh represented by the matrix elements $O_{qp}$, the transfer function to each output of the programmable spectral filter is given by Eq. 9 where $x(\Delta\omega)$ is the input spectrum to the PIC and $y_{q}(\Delta\omega)$ is the spectrum at the $q^{th}$ output.
@@
$y_{q}(\Delta\omega)=[\sum_{p=0}^{N-1}a_{p}O_{qp}e^{-j(k_{p}\Delta\omega\Delta{t}\ -2\pi\gamma)}]x(\Delta\omega)$ (11)
@@
An additional dynamic, wavelength-independent phase profile may be introduced using programmable phase shifters to perturb the static phase profile generated by the array of waveguide delay lines. This dynamic phase profile may be generated either using the $\phi_{00}$, $\phi_{01}$, and $\phi_{02}$ phase shifters in the $m=0$ layer of the photonic mesh or the $\phi_{0}$, $\phi_{1}$, and $\phi_{2}$ phase shifters of the power splitting sub-circuit. Whether this phase profile is introduced prior to or following the array waveguide delay lines is of no consequence. By programming these phase shifters to introduce a linear tilt in addition to the phase shifts determined by self-configuration, the center wavelength of a preconfigured filter function may be shifted. The programmable linear phase profile is defined by the terms $e^{-j2\pi\gamma{p}}$. By adjusting the value of gamma between 0 and 1, the center wavelength of the preconfigured filter function may be tuned continuously across an FSR. This principle is demonstrated for a $N=4,M=1$ spectral filter for a selection of $\gamma$ values in figure 4. The ability to modulate the center wavelength of a preconfigured filter function is a prerequisite for applications such as correlation spectroscopy used in environmental and biomedical sensing [15].
@@
It can further be observed from Fig 3 that outputs $q=1,2,$ and $3$ have multiple transmission peaks with varying relative amplitudes and frequency spacings. The relative amplitude and frequency spacing are functions of the phase shifter settings of the remaining layers of the photonic mesh. These parameters may be adjusted dynamically and arbitrarily, enabling complex functionalities such as intra-FSR pulse shaping. Were this device to act as the wavelength selective component of an external cavity laser, these multiple transmission peaks may be tuned to enable multimode lasing within a single FSR.
@@ Discussion
Among potential applications, this device enables tunable pump wavelength rejection in Raman spectroscopy, continuously tunable bandpass filters for use in external cavity lasers, tunable multimode operation for external cavity lasers, and reprogrammable synthetic spectra generation for use in correlation spectroscopy. The ability to modulate the central wavelength of a preconfigured filter function, enabled by this device, is a prerequisite of filters used in correlation spectroscopy. Further, we have demonstrated the capability of the device to perform DWDM with inter-channel crosstalk between -25 dB and -40 dB. Unlike typical DWDM systems that operate on a static wavelength grid, mapping individual wavelengths sequentially to individual output channels, this device may dynamically shift the wavelength grid and dynamically remap any input wavelength channel to any output waveguides. In essence, this further introduces the functionality of a switching network.
@@
Materials and Methods
[…]
The single-mode waveguide used throughout the architecture has a cross-section that is 220 nm tall by 500 nm wide. The fundamental quasi-TE mode has an effective and group index of $n_{r}=2.459$ and $n_{g}=4.056$, respectively, at a central wavelength of 1550 nm. These values have been simulated using the open-source Eigenmode and FDTD solver EMopt. The coupling of light, both on and off chip, has been managed using surface grating couplers designed to convert between these single-mode silicon waveguides and cleaved single-mode fibers (SMF28). The input and output grating couplers are designed to emit at approximately $12^{\circ}$ from vertical with approximately 6 dB of insertion loss per coupler.
We use a HP 81680A tunable laser as an input source to define the pilot frequencies throughout this work. The output of the laser is fiber coupled to a Thorlabs FPC562 manual polarization controller to manage the input polarization state. The output fiber of the polarization controller has been cleaved to ensure good mode matching with surface grating couplers. Both input and output fibers are precisely aligned to their respective grating couplers with a Thorlabs Nanomax 600 series 6-axis stage.
An over-head observation microscope has been assembled to aid coarse alignment between the input/output fibers and their respective grating couplers. A Xenics Bobcat 640 In-GaAs camera is placed in the detector plane of the microscope, enabling in-situ observation of the incident and output fields. This same microscope assembly enables detection of power at the monitor taps used for calibration of the power splitting subcircuit. Fine alignment between input and output fibers is optimized based on throughput power detected at a HP 81531A photodetector. This same detector is used to measure the output spectra of the programmable filter.
[…]
An interface PCB has been designed to route electrical driving signals from the NI DAC to the PIC. A third-party vendor, Silitronics, has been used for the wire-bonding between the PCB bond pads and the PIC bond pads. During the wire-bonding process, the PIC is permanently fixed to the PCB with a thermally conductive adhesive. For thermal management during the operation of the spectrometer, the assembly is secured to an aluminum mount for heat sinking. Additionally, a pocket has been milled into the aluminum mount for a thermo-electric cooling (TEC) unit to be inserted directly beneath the PIC.
[…]
Acknowledgments This work has been funded by the Air Force Office of Scientific Research (FA9550-21-1-0312, FA9550-23-1-0307), Ames Research Center (80NSSC24M0033), Marie Skłodowska-Curie Actions (101067268), and Stanford Engineering.
[…]
Author Details 1Stanford University, Ginzton Laboratory, 348 Via Pueblo Mall, Stanford CA 94305. 2Department of Physics and Technology, UiT The Arctic University of Norway, NO-9037 Tromsø
[…]
Author Contributions D.A.B.M, S.F. and O.S. supervised the project. D.A.B.M, S.F., O.S. M.V., A.K. C.C. and C.V developed the analytical model. A.K., M.V., and C.V. designed the PIC. C.V and A.K. built the setups. M.V., A.K.,and C.V. performed experimental characterization. C.V wrote the manuscript with input from all authors.
[…]
Data availability Data underlying the results presented in this paper is available from the corresponding authors upon request.
[…]
Disclosures The authors declare no conflicts of interest.